# Sample Compression Unleashed: New Generalization Bounds for Real Valued Losses

### v2

@@ 4 Experiments
In this section, we show the versatility of our results by training different models using the P2L algorithm. [^1] Our code is available at https://github.com/GRAAL-Research/pick-to-learn. In Section 4.1, we train neural networks on binary classification problems and compare our new results to the pre-existing sample compression results. We empirically validate that our bounds are almost as tight as the binomial bound, all the while not suffering from the numerical optimization problem of Theorem 1 and being defined in the inconsistent case, where the P2L bound of Theorem 2 is undefined. In Section 4.2, we train CNNs on the MNIST dataset and present generalization bounds on the (bounded) cross-entropy loss. As no previous sample-compression bound is defined for real-valued losses, we compare our result to a PAC-Bayesian theorem. In Section 4.3, we use P2L to train decision forests on regression datasets and give generalization bounds on the root mean squared error (RMSE), an unbounded loss function, under the assumption that it is sub-Gaussian. Finally, in Section 4.4, we fine-tune DistilBERT, a 66M parameters language model, on a review polarity classification problem. We obtain tight bounds simultaneously on the zero-one loss and the cross-entropy loss, demonstrating that our new theorem is independent of the number of parameters of the model.
@@
Each experiment is run five times with different seeds (with the exception of amazon polarity). In all tables, we present the mean and standard deviation of the metrics over five seeds. The datasets are separated into three parts: the training, validation and test set. The validation set is built using 10% of the training set. If the dataset doesn’t have a built-in test set, we build a test set first, using 10% of the dataset, and then build the validation set. When computing the bounds, we use $\delta=0.01$. All baselines are trained on the whole dataset using stochastic gradient descent for 200 epochs or until the model achieves zero errors on the training set. The random forests (Breiman,, 2001) are trained on the whole dataset, with no modifications to the training algorithm. All the hyperparameters for all the experiments can be found in Appendix A.
4.1 Binary MNIST
Table 1: Results for the CNNs trained using P2L on the binary MNIST problems. The results displayed obtained the tightest P2L bound. All metrics presented are in percent (%), with the exception of $\operatorname{|\mathbf{i}|}$.
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound P2L bound $\operatorname{|\mathbf{i}|}$ Baseline test error
MNIST08 0.33$\pm$0.17 0.25$\pm$0.10 5.05$\pm$0.16 5.00$\pm$0.16 1.04$\pm$0.04 92.0$\pm$3.6 0.22$\pm$0.05
MNIST17 0.20$\pm$0.08 0.38$\pm$0.16 4.33$\pm$0.21 4.29$\pm$0.21 0.86$\pm$0.05 84.0$\pm$5.2 0.17$\pm$0.03
MNIST23 0.39$\pm$0.12 0.27$\pm$0.10 8.20$\pm$0.34 8.15$\pm$0.34 1.86$\pm$0.09 175.6$\pm$9.5 0.16$\pm$0.05
MNIST49 0.82$\pm$0.11 0.77$\pm$0.17 10.52$\pm$0.37 10.47$\pm$0.37 2.53$\pm$0.11 237.0$\pm$11.0 0.44$\pm$0.07
MNIST56 0.46$\pm$0.12 0.47$\pm$0.15 6.29$\pm$0.22 6.24$\pm$0.22 1.35$\pm$0.06 117.0$\pm$5.2 0.30$\pm$0.08
Table 2: Results for the CNNs trained using P2L on the binary MNIST problems and stopped at the iteration with the minimum $\mathrm{kl}$ bound. The results displayed obtained the tightest $\mathrm{kl}$ bound. Metrics are in percents (%), except $\operatorname{|\mathbf{i}|}$.
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound Train error $\operatorname{|\mathbf{i}|}$ Baseline test error
MNIST08 0.49$\pm$0.39 0.49$\pm$0.26 4.71$\pm$0.25 5.33$\pm$0.62 0.24$\pm$0.23 66.0$\pm$15.0 0.22$\pm$0.05
MNIST17 0.45$\pm$0.18 0.48$\pm$0.11 3.70$\pm$0.21 4.37$\pm$0.11 0.23$\pm$0.08 50.0$\pm$8.9 0.17$\pm$0.03
MNIST23 0.74$\pm$0.28 0.84$\pm$0.21 6.56$\pm$0.38 8.09$\pm$0.64 0.64$\pm$0.32 84.0$\pm$21.5 0.16$\pm$0.05
MNIST49 1.16$\pm$0.31 1.13$\pm$0.24 8.60$\pm$0.46 9.61$\pm$0.68 0.51$\pm$0.28 134.0$\pm$24.2 0.44$\pm$0.07
MNIST56 0.94$\pm$0.09 0.70$\pm$0.20 5.42$\pm$0.31 6.49$\pm$0.81 0.43$\pm$0.23 66.0$\pm$10.2 0.30$\pm$0.08
We create binary classification datasets by choosing two digits from the MNIST dataset (LeCun et al.,, 1998), e.g., choosing all the datapoints labeled 00 and $8$ to build the dataset MNIST08. We create five datasets: MNIST08, MNIST17, MNIST23, MNIST49 and MNIST56. Starting from randomly initialized neural networks, we train a MLP and a CNN using P2L on each dataset. More details are given in Section A.1.1.
We present our results for the CNN in Table 1. All the results for the MLP can be found in Section A.1.1. The error on the training set is zero for all predictors returned by P2L. The results presented achieved the tightest P2L bound for each dataset. For reference, the reported “baseline test error” corresponds to the results of the best baseline model based on validation error. For both architectures, using P2L only incurs a slight increase of the test error compared to the baseline, whilst the model is trained on a very small percentage of the dataset, ranging from 0.7% to 3.4%. Finally, even though the P2L bound is much tighter than the proposed $\mathrm{kl}$ bound, our result is much more general, as it holds for any real-valued loss functions and in the non-consistent case. Moreover, our bounds hold uniformly over all iterations of the models trained using P2L. After training, one can use any checkpoint of the model and still obtain a valid bound, which gives control over a trade-off between the training error, the generalization bound and the validation error. In Fig. 1, we present the behavior of the bound throughout the P2L iterations. The minimal $\mathrm{kl}$ bound happens at about half the final number of iterations, leading to a smaller compression set and a tighter bound, as also reported in Table 2. In comparison to the previous results, the test error is about twice as high as the test error of the fully trained model (Table 1). However, the models were trained on very small portions of the dataset, with the model on MNIST17 being trained on 0.42% of the dataset and still achieving a test error of 0.48%. Finally, we observe that, in this setting, our new $\mathrm{kl}$ bound is much tighter than the binomial approximation of Laviolette et al., (2005).
@@ 4.2 MNIST
Model Train loss Test loss $\mathrm{kl}$ bound $\operatorname{|\mathbf{i}|}$ Baseline test loss
P2L 0.0008$\pm$0.0006 0.0480$\pm$0.0073 0.7142$\pm$0.1773 275.20$\pm$82.46 0.0499$\pm$0.0108
PBB 0.0092$\pm$0.0005 0.0045$\pm$0.0004 0.0112$\pm$0.0005 -
We now train convolutional neural networks composed of two convolutional layers and two fully connected layers. We pre-train the model using stochastic gradient descent on a subset of the dataset and then use P2L to fine-tune the model on the train set. The size of the pre-training subset is an hyperparameter. We use the same training setting as in Section 4.1 and use the extension of P2L that adds multiple datapoints to the compression set at a time, with batch size $R=32$, as defined by Algorithm 2 of Paccagnan et al., (2024). For comparison, we also train probabilistic neural networks (PNN) using the PAC-Bayes with Backprop (PBB) approach of Pérez-Ortiz et al., (2021). They train the model by minimizing the PAC-Bayesian $\mathrm{kl}$ bound of Theorem 10. See Section A.1.2 for details.
@@
In Table 3, we report the bound values for the bounded cross-entropy loss (see Appendix for classification error). We observe that the PBB algorithm gives a tighter generalization bound than the one of P2L. This gap can be explained by the fact that PBB jointly optimizes the train error and the $\mathrm{KL}$ divergence, whilst we have almost no control on the minimization of the bound. Indeed, the heuristic of the P2L algorithm, which is to choose the datapoints over which the model incurs the greatest losses, doesn’t give control on the trade-off between the decrease of the error and the increase of the complexity term. Moreover, for a large dataset, the binomial coefficient increases rapidly when the compression set size increases. However, using our bounds with the P2L algorithm has multiple advantages over the PBB algorithm. First of all, PBB needs to train twice as many parameters, as it fits both the mean and standard deviation of the distributions over the parameters. Secondly, computing the PAC-Bayesian bound necessitates a step of Monte Carlo sampling to determine the average error of the model. For 5000 steps of Monte Carlo sampling, the error over the dataset will be computed 5000 times, instead of only once with P2L. Finally, our bound doesn’t take into account the number of parameters of the model, whilst the $\mathrm{KL}$ divergence in Theorem 10 is a sum of the $\mathrm{KL}$ divergence of the distribution of each parameter of the model.
4.3 Regression forests
Table 4: Results for the decision forests trained using P2L. We report the RMSE achieved by the models and the generalization bounds on the RMSE.
Dataset Train loss Validation loss Test loss $\mathrm{kl}$ bound Linear bound $\operatorname{|\mathbf{i}|}$ Baseline test loss $\ell^{\max}$
Powerplant 5.23$\pm$2.23 5.23$\pm$2.18 5.37$\pm$2.33 11.08$\pm$5.04 12.79$\pm$5.91 29.20$\pm$17.81 3.59$\pm$0.13 90.6
Infrared 0.27$\pm$0.03 0.29$\pm$0.04 0.30$\pm$0.03 1.08$\pm$0.08 1.16$\pm$0.08 19.20$\pm$5.49 0.23$\pm$0.01 4.26
Airfoil 3.57$\pm$0.34 3.91$\pm$0.21 3.88$\pm$0.39 14.78$\pm$1.39 14.84$\pm$1.26 46.80$\pm$15.93 2.10$\pm$0.15 45.13
Parkinson 7.59$\pm$0.44 7.75$\pm$0.50 7.73$\pm$0.36 12.13$\pm$0.37 11.98$\pm$0.41 22.60$\pm$10.59 2.23$\pm$0.16 41.37
Concrete 8.59$\pm$1.10 8.74$\pm$0.79 8.78$\pm$1.07 30.18$\pm$1.61 31.15$\pm$1.43 26.20$\pm$7.28 4.70$\pm$0.36 90.63
@@ 4.3 Regression forests
Table 5: Results for the amazon polarity dataset. The results displayed for P2L obtained the lowest $\mathrm{kl}$ bound on the error, whilst the baseline was chosen by the lowest validation error.
Model Error (%) Cross-entropy loss
Train Validation Test $\mathrm{kl}$ bound Binomial bound Train Validation Test $\mathrm{kl}$ bound $\operatorname{|\mathbf{i}|}$
P2L 4.25 5.24 5.37 14.67 21.63 0.1180 0.1465 0.1486 0.9992 1472
Baseline 3.13 4.07 4.19 - - 0.0902 0.1151 0.1156 - -
@@
In order to show the wide applicability of our bounds, we train decision forests on regression problems: Powerplant (Tüfekci,, 2014), Infrared (Wang et al.,, 2021), Airfoil (Brooks et al., 1989b, ), Parkinson (Tsanas et al.,, 2009) and Concrete (Yeh, 1998b, ). These datasets range from a training set size of 827 to 7751 and range from a number of features of 4 to 33. To the best of our knowledge, no sample compression bounds exist for this setting. We adapt the P2L algorithm to this regression problem (see Algorithm 2 in appendix), which differs from the original one, designed only for classification problems where zero training error is achievable (consistent case). At each P2L iteration, we add a single datapoint to the compression set in order to train the random forest. The selected datapoint is the one with the largest root mean squared error (RMSE). Then, the trees are retrained completely on the compression set. As the minimal RMSE that can be achieved is dependent on the dataset, setting a predetermined threshold is not a suitable stopping criterion. Thus, we train the model until the validation loss has not decreased for a number of iterations. To compute the bounds, we need the loss to be either bounded or sub-Gaussian. As tree-based models predict the mean of the targets of each datapoint assigned to a leaf, their outputs are bounded by the extrema of the data. Thus, if we assume that the target space is bounded, the loss will be bounded. To compute the $\mathrm{kl}$ bound, we assume that the target space is bounded. The maximum value of the loss $\ell^{\max}$ on each dataset is reported in Table 4. To compute the linear bound, we assume that the loss is sub-Gaussian. We discuss in more details these assumptions and the way of defining the extrema in Section A.1.3.
@@
We present the results in Table 4. The models are selected based on the smallest $\mathrm{kl}$ bound. We observe that the models trained with P2L are able to obtain competitive results with respect to the test error of our baseline, random forests trained on the whole dataset. We report these results in the column “baseline test loss” of Table 4, where the models were chosen by their validation loss. As the value of the bounds is always much smaller than $\ell^{\max}$, we can observe that our bounds are tight and non-vacuous. The generalization guarantees given by the bound using the linear function are competitive to the $\mathrm{kl}$ and are even tighter on the Parkinson dataset.
Experiments with regression trees can be found in Section A.1.3. Training trees using P2L lead to underfitted trees that were not competitive w.r.t. the baseline (see Table 11). To the best of our knowledge, these results are the first generalization bounds for regression trees.

### v3

@@ 4 EXPERIMENTS
In this section, we show the versatility of our results by training different models using the P2L algorithm. [^3] Our code is available at https://github.com/GRAAL-Research/pick-to-learn. In Section 4.1, we train neural networks on binary classification problems and compare our new results to the pre-existing sample compression ones. We empirically validate that our bounds are almost as tight as the binomial bound, all the while not suffering from the numerical optimization problem of Theorem 1 and being defined in the inconsistent case, where the P2L bound of Theorem 2 is undefined. In Section 4.2, we train CNNs on the MNIST dataset and present generalization bounds on the (bounded) cross-entropy loss. As no previous sample-compression bound is defined for real-valued losses, we compare our result to a PAC-Bayesian theorem. In Section 4.3, we use P2L to train tree-based models on regression datasets and give generalization bounds on the root mean squared error (RMSE), an unbounded loss function, under the assumption that it is sub-Gaussian. Finally, in Section 4.4, we fine-tune DistilBERT, a 66M parameters language model, on a review polarity classification problem. We obtain tight bounds simultaneously on the zero-one loss and the cross-entropy loss, demonstrating that our new theorem is independent of the number of parameters of the model.
@@
Each experiment is run five times with different seeds. We report the mean and standard deviation of the metrics over these five repetitions. The datasets are separated into three parts: the training, validation and test set. When a dataset doesn’t have a built-in test set, we create it using 10% of the samples. Of the remaining samples, 10% are used for the validation set size and 90% for the training set. When computing the bounds, we use $\delta=0.01$. The hyperparameters used for the experiments can be found in Appendix A.
@@
Learning algorithm Train loss Test loss $\mathrm{kl}$ bound $\operatorname{|\mathbf{i}|}/n$ Baseline test loss
P2L 0.0014$\pm$0.0006 0.0467$\pm$0.0056 0.8279$\pm$0.0256 1.13$\pm$0.24 0.0499$\pm$0.0042
PBB 0.0092$\pm$0.0005 0.0045$\pm$0.0004 0.0112$\pm$0.0005 -
4.1 Binary MNIST
We create binary classification datasets by extracting pairs of digits from the MNIST dataset (LeCun et al.,, 1998), e.g., selecting the datapoints labeled 00 and $8$ to build the dataset MNIST08. We create five datasets: MNIST08, MNIST17, MNIST23, MNIST49 and MNIST56. Starting from randomly initialized neural networks, we train a MLP and a CNN using P2L on each dataset. More details are given in Section A.1.1.
Table 1 presents results for the CNN in the consistent case. The results for the MLP can be found in Section A.1.1. The error on the training set is zero for all predictors returned by P2L. The presented results achieve the tightest P2L bound for each dataset. The reported “baseline test error” corresponds to the results of a standard neural network optimized on the whole training set by stochastic gradient descent for 200 epochs or until the model achieves zero training errors; the selected hyperparmeters are the ones minimizing the validation error. For both CNN and MLP architectures, using P2L only incurs a slight increase of the test error compared to the baseline, whilst the model is trained on a fraction of the dataset, ranging from 0.7% to 3.4%. Finally, even though the P2L bound is much tighter than the proposed $\mathrm{kl}$ bound, our result is much more general, as it holds for any real-valued loss functions and in the non-consistent case. Moreover, our bounds hold uniformly over all iterations of the models trained using P2L. After training, one can use any checkpoint of the model and still obtain a valid bound, which gives control over a trade-off between the training error, the generalization bound and the validation error.
As for the inconsistent case, Figure 1 presents the behavior of the bound throughout the P2L iterations. We observe that the minimal $\mathrm{kl}$ bound value happens at about half the final number of iterations, leading to a smaller compression set and a tighter bound, as also reported in Table 2. Recall that the P2L bound cannot be computed in this case, as the model do not reach zero errors. In comparison to the previous consistent results (Table 1), the test error of Table 2 are about twice as high as the fully trained model. However, the inconsistent models are trained on very small portions of the dataset, with the MNIST17 model being trained on 0.42% of the dataset and still achieving a test error of 0.48%. Finally, we observe that, in this setting, our new $\mathrm{kl}$ bound is much tighter than the binomial approximation of Laviolette et al., (2005).
@@ 4.2 MNIST
Table 4: Results for the decision forests trained using P2L. We report the RMSE achieved by the models with the smallest $\mathrm{kl}$ bound. The ratio $\operatorname{|\mathbf{i}|}/n$ is presented in percents (%).
Dataset Train loss Validation loss Test loss $\mathrm{kl}$ bound Linear bound $\operatorname{|\mathbf{i}|}/n$ Baseline test loss $\ell^{\max}$
Powerplant 6.11$\pm$0.89 6.23$\pm$0.70 6.31$\pm$0.95 13.69$\pm$0.27 15.92$\pm$0.47 0.51$\pm$0.17 3.59$\pm$0.13 90.6
Infrared 0.27$\pm$0.03 0.29$\pm$0.04 0.30$\pm$0.03 1.08$\pm$0.08 1.16$\pm$0.08 2.32$\pm$0.66 0.23$\pm$0.01 4.26
Airfoil 3.67$\pm$0.16 4.03$\pm$0.37 3.90$\pm$0.18 14.19$\pm$0.49 14.25$\pm$0.42 3.23$\pm$0.46 2.10$\pm$0.15 45.13
Parkinson 7.79$\pm$0.33 7.79$\pm$0.28 7.84$\pm$0.27 12.24$\pm$0.27 12.02$\pm$0.23 0.43$\pm$0.10 2.23$\pm$0.16 41.37
Concrete 8.18$\pm$0.91 8.70$\pm$1.00 8.48$\pm$1.41 31.68$\pm$1.63 32.49$\pm$1.52 3.81$\pm$0.82 4.70$\pm$0.36 90.63
Table 5: Results for the decision trees trained using P2L. We report the RMSE achieved by the models with the smallest $\mathrm{kl}$ bound. The ratio $\operatorname{|\mathbf{i}|}/n$ is presented in percents (%)
Dataset Train loss Validation loss Test loss $\mathrm{kl}$ bound Linear bound $\operatorname{|\mathbf{i}|}/n$ Baseline test loss $\ell^{\max}$
Powerplant 11.66$\pm$3.00 11.83$\pm$3.12 12.00$\pm$3.04 23.40$\pm$1.59 24.15$\pm$1.54 0.94$\pm$0.49 4.07$\pm$0.13 90.6
Infrared 0.48$\pm$0.10 0.47$\pm$0.11 0.48$\pm$0.07 1.34$\pm$0.03 1.33$\pm$0.06 2.10$\pm$1.06 0.27$\pm$0.03 4.26
Airfoil 11.02$\pm$1.71 10.89$\pm$1.38 11.10$\pm$1.93 18.87$\pm$1.97 18.14$\pm$1.75 1.00$\pm$0.03 3.01$\pm$0.19 45.13
Parkinson 13.93$\pm$2.83 13.84$\pm$2.76 13.99$\pm$2.92 17.75$\pm$3.34 15.10$\pm$7.23 0.25$\pm$0.00 3.20$\pm$0.15 41.37
Concrete 26.08$\pm$4.96 25.05$\pm$4.38 26.40$\pm$4.05 45.79$\pm$4.26 44.47$\pm$4.24 1.51$\pm$0.33 6.22$\pm$0.91 90.63
We now train convolutional neural networks composed of two convolutional layers and two fully connected layers. We pre-train the model using stochastic gradient descent on a subset of the dataset and then use P2L to fine-tune the model on the train set. The size of the pre-training subset is an hyperparameter. We use the same training setting as in Section 4.1 and use the extension of P2L that adds multiple datapoints to the compression set at a time, with batch size $R=32$, as defined by Algorithm 2 of Paccagnan et al., (2024). For comparison, we also train probabilistic neural networks (PNN) using the PAC-Bayes with Backprop (PBB) approach of Pérez-Ortiz et al., (2021), which optimize a PAC-Bayesian $\mathrm{kl}$ bound (Theorem 11 in appendix).
@@
Table 3 reports the bound values for the bounded cross-entropy loss (see Table 11 for classification error). We observe that the PBB algorithm gives a tighter generalization bound than the one of P2L. This gap can be explained by the fact that PBB jointly optimizes the train error and the $\mathrm{KL}$ divergence, whilst we have almost no control on the minimization of the bound. Indeed, the heuristic of the P2L algorithm, which is to choose the datapoints over which the model incurs the greatest losses, doesn’t give control on the trade-off between the decrease of the error and the increase of the complexity term. Moreover, for a large dataset, the binomial coefficient increases rapidly when the compression set size increases. However, using our bounds with the P2L algorithm has multiple advantages over the PBB algorithm. First of all, PBB needs to train twice as many parameters, as it fits both the mean and standard deviation of the distributions over the parameters. Secondly, computing the PAC-Bayesian bound necessitates a step of Monte Carlo sampling to determine the average error of the model. For 5000 steps of Monte Carlo sampling, the prediction over the dataset is computed 5000 times, instead of only once with P2L. Finally, our bound doesn’t take into account the number of parameters of the model, whilst the $\mathrm{KL}$ divergence in Theorem 11 is a sum of the $\mathrm{KL}$ divergence of the distribution of each parameter of the model.
4.3 Regression with tree-based models
@@
Table 6: Results for the amazon polarity dataset. The results displayed for P2L obtained the lowest $\mathrm{kl}$ bound on the error, whilst the baseline was chosen by the lowest validation error. The ratio $\operatorname{|\mathbf{i}|}/n$ is presented in percents (%).
Learning algorithm Error (%) Cross-entropy loss
Train Test $\mathrm{kl}$ bound Binomial bound Train Test $\mathrm{kl}$ bound $\operatorname{|\mathbf{i}|}/n$
P2L 4.73$\pm$1.09 5.60$\pm$1.19 13.91$\pm$2.73 21.85$\pm$3.29 0.1199$\pm$0.0118 0.1478$\pm$0.0182 0.8594$\pm$0.1622 0.79$\pm$0.23
Baseline 3.11$\pm$0.02 4.19$\pm$0.00 - - 0.0912$\pm$0.0010 0.1158$\pm$0.0002 - -
@@
In order to show the wide applicability of our bounds, we train decision forests on regression problems: Powerplant (Tüfekci,, 2014), Infrared (Wang et al.,, 2021), Airfoil (Brooks et al., 1989b, ), Parkinson (Tsanas et al.,, 2009) and Concrete (Yeh, 1998b, ). These datasets range from a training set size of 827 to 7751 and range from a number of features of 4 to 33. To the best of our knowledge, no sample compression bounds exist for this setting. Previous results were presented for linear regression under an $\ell_{p}$ loss (Attias et al.,, 2018, 2023, 2024) or for boosting real-valued learner in a binary classification setting (Hanneke et al.,, 2019), none of which are equivalent to our setting. We adapt the P2L algorithm to this regression problem (see Algorithm 2 in appendix), which differs from the original one, designed only for classification problems where zero training error is achievable (consistent case). At each P2L iteration, we add a single datapoint to the compression set in order to train the forest. The selected datapoint is the one with the largest root mean squared error (RMSE). Then, the trees are retrained from scratch on the compression set. As the minimal RMSE that can be achieved is dependent on the dataset, setting a predetermined threshold is not a suitable stopping criterion. Thus, we train the model until the validation loss has not decreased for a given number of iterations. To compute the bounds, we need the loss to be either bounded or sub-Gaussian. As tree-based models predict the mean of the targets of each datapoint assigned to a leaf, their outputs are bounded by the extrema of the data. To compute the $\mathrm{kl}$ bound, we assume that the target space is bounded by the maximum value of the loss $\ell^{\max}$ reported in Table 4. To compute the linear bound, we assume that the loss is sub-Gaussian. We discuss in more details these assumptions and the way of defining the extrema in Section A.1.3.
@@
Table 4 contains the results of the models selected based on the smallest $\mathrm{kl}$ bounds. We observe that the models trained with P2L are able to obtain competitive results with respect to the test error of standard random forests trained on the whole dataset. We report latter results in the column “baseline test loss”, where the models are chosen by their validation loss. As the values of the bounds are much smaller than $\ell^{\max}$, we conclude that our bounds are tight and non-vacuous. The generalization guarantees given by the bounds relying on the linear function are competitive to the ones relying on the $\mathrm{kl}$; they are even tighter on the Parkinson dataset.
Following a similar experiment setting, we trained regression trees with P2L. As we can see in Table 5, training trees using P2L leads to underfitted models that are not competitive with respect to the baseline. Indeed, as the trees are trained on a small subset of the data, they are restricted to be less complex than trees trained on the whole dataset. When selecting the models by the smallest validation loss (see Table 13), we observe that the models achieve better performance, as the compression sets are much larger, but also suffer from worsened bounds.